### v1

# Data Bias According to Bipol: Men are Naturally Right and It is the Role of Women to Follow Their Lead

Abstract
We introduce new large labeled datasets on bias in 3 languages and show in experiments that bias exists in all 10 datasets of 5 languages evaluated, including benchmark datasets on the English GLUE/SuperGLUE leaderboards. The 3 new languages give a total of almost 6 million labeled samples and we benchmark on these datasets using SotA multilingual pretrained models: mT5 and mBERT. The challenge of social bias, based on prejudice, is ubiquitous, as recent events with AI and large language models (LLMs)have shown. Motivated by this challenge, we set out to estimate bias in multiple datasets. We compare some recent bias metrics and use bipol, which has explainability in the metric. We also confirm the unverified assumption that bias exists in toxic comments by randomly sampling 200 samples from a toxic dataset population using the confidence level of 95% and error margin of 7%. Thirty gold samples were randomly distributed in the 200 samples to secure the quality of the annotation. Our findings confirm that many of the datasets have male bias (prejudice against women), besides other types of bias. We publicly release our new datasets, lexica, models, and codes.
Caution: This paper contains examples of what some may consider as stereotypes or offensive text.
1 Introduction
The problem of social bias in data is a pressing one. Recent news about social bias of artificial intelligence (AI)systems, such as Alexa111bbc.com/news/technology-66508514 and ChatGPT,222bloomberg.com/news/newsletters/2022-12-08/chatgpt-open-ai-s-chatbot-is-spitting-out-biased-sexist-results shows that the age-old problem persists with data, which is used to train machine learning (ML)models. Social bias is the inclination or prejudice for, or against, a person, group or idea, especially in a way that is considered to be unfair, which may be based on race, religion or other factors (Bellamy et al., 2018; Antoniak & Mimno, 2021; Mehrabi et al., 2021; Alkhaled et al., 2023). It can also involve stereotypes that generalize behavior to groups (Brownstein, 2019). It can unfairly skew the output of ML models (Klare et al., 2012; Raji et al., 2020). Languages with fewer resources than English are also affected (Rescigno et al., 2020; Chávez Mulsa & Spanakis, 2020; Kurpicz-Briki, 2020). For example, in Italian, the female gender is under-represented due to the phenomena such as the “inclusive masculine" (when the masculine is over-extended to denote groups of both male and female referents) (Luccioli et al., ; Vanmassenhove & Monti, 2021).
In this work, we are motivated to address the research question of how much bias exists in the text data of multiple languages, if at all bias exists in them. We particularly investigate 6 benchmark datasets on the English GLUE/SuperGLUE leaderboards (Wang et al., 2018; 2019) and one dataset each for the other 4 languages: Italian, Dutch, German, and Swedish. First, we train SotA multilingual Text-to-Text Transfer Transformer (mT5)(Xue et al., 2021) and multilingual Bidirectional Encoder Representations from Transformers (mBERT) models for bias classification on the multi-axes bias dataset (MAB)(Alkhaled et al., 2023) for each language. For the evaluations, we search through the literature to compare different metrics or evaluation methods as shown in Table 1 and discussed in Section 2. This motivates our choice of bipol, the multi-axes bias metric, which we then compare in experiments with a lexica baseline method. In addition, to confirm the unverified assumption that toxic comments contain bias (Sap et al., 2020; Alkhaled et al., 2023), we annotate 200 randomly-selected samples from the training set of the English MAB.
Metric/Evaluator Axis Terms
Winogender (Rudinger et al., 2018) 1 60
WinoBias (Zhao et al., 2018) 1 40
StereoSet (Nadeem et al., 2021) 4 321
GenBiT (Sengupta et al., 2021) 1 -
CrowS-Pairs (Nangia et al., 2020) 9 3,016
Bipol (Alkhaled et al., 2023) $>$2, $<$13 $>$45, $<$466
Table 1: Comparison of some bias evaluation methods.
Our Contributions
  • We make available new large labeled datasets on bias of almost 2 million samples each for 3 languages: Italian, Dutch, and German.333github.com/LTU-Machine-Learning/bipolmulti
  • We make available lexica of sensitive terms for bias detection in the 3 languages.
  • We confirm the unverified assumption in the underlying datasets of MAB (Social Bias Inference Corpus v2 (SBICv2)and Jigsaw) (Alkhaled et al., 2023) that toxic comments contain bias.
The rest of this paper is organized as follows. In Section 2, we discuss the literature review of related work. In Section 3, we briefly discuss the bipol metric. In Section 4, we explain the steps involved in the methodology and the datasets we use. In Section 5, we present our findings and discuss them. In Section 6, we end with the conclusion and possible future work.
2 Literature Review
Although English usually gets more support and attention in the literature, there have been attempts at measuring and mitigating bias in other languages. Testing for the presence of bias in Italian often has a contrastive perspective with English, with a focus on gender bias (Gaido et al., 2021; Rescigno et al., 2020). MuST-SHE (Bentivogli et al., 2020) and gENder-IT (Vanmassenhove & Monti, 2021) are examples of gender bias evaluation sets. Going beyond gender bias, Kurpicz-Briki & Leoni (2021) and Huang et al. (2020) also identified biases related to people’s origin and speakers’ age. It is essential to remember that the mentioned biases can be vehicles for misogynous and hateful discourse (El Abassi & Nisioi, 2020; Attanasio et al., 2022; Merenda et al., 2018).
Bias studies for Dutch mostly consider binary gender bias. Chávez Mulsa & Spanakis (2020) investigate gender bias in Dutch static and contextualized word embeddings by creating Dutch versions of the Word/Sentence Embedding Association Test (WEAT/SEAT) (Caliskan et al., 2017; May et al., 2019). WEAT measures bias in word embeddings and can be limited in scope, in addition to having sensitivity to seed words. McCurdy & Serbetci (2020) perform a similar evaluation in a multilingual setup to compare the effect of grammatical gender saliency across languages. Several works use different NLP techniques to evaluate bias in corpora of Dutch news articles (Wevers, 2019; Kroon et al., 2020; Kroon & van der Meer, 2021; Fokkens et al., 2018) and literary texts (Koolen & van Cranenburgh, 2017).
In Kurpicz-Briki (2020), bias is measured with regards to place of origin and gender in German word embeddings using WEAT (Caliskan et al., 2017). In Kurpicz-Briki & Leoni (2021), an automatic bias detection method (BiasWords) is presented, through which new biased word sets can be identified by exploring the vector space around the well-known word sets that show bias. In the template-based study of Cho et al. (2021), on gender bias in translations, the accuracy of gender inference was measured for multiple languages including German. It was found that, particularly for German, the inference accuracy and disparate impact were lower for female than male, implying that certain translations were wrongly performed for cases that required female inference. Since German is a grammatically gendered, morphologically rich language, Gonen & Goldberg (2019) found that debiasing methods of Bolukbasi et al. (2016) were ineffective on german word embeddings.
For Swedish, the main focus of bias research appears to be on gender. Sahlgren & Olsson (2019) show with their experiments that gender bias is present in pretrained Swedish language models. Katsarou et al. (2022) and Precenth (2019) found that the male gender tends to be associated with higher-status professions. A study with data from mainstream news corpora by Devinney et al. (2020) shows that women are associated with concepts like family, communication and relationships.
3 Bipol
The bipol metric uses a two-step mechanism for estimating bias in text data: binary classification and sensitive term evaluation using lexica. It has maximum and minimum values of 1 and 0, respectively. Bipol is expressed in Equations 1b and 1c from the main Equation 1a, where $b_{c}$ is the classification component and $b_{s}$ is the sensitive term evaluation component. Details about the metric are discussed in Alkhaled et al. (2023).
$\mathit{b}=\begin{cases}\mathit{b_{c}}.\mathit{b_{s}},&\text{if $b_{s}>0$}\\ \mathit{b_{c}},&\text{otherwise}\end{cases}$ (1a)
$\mathit{b_{c}}=\frac{tp+fp}{tp+fp+tn+fn}$ (1b)
$\mathit{b_{s}}=\frac{1}{r}\sum_{t=1}^{r}{\left(\frac{1}{q}\sum_{x=1}^{q}{\left(\frac{|\sum_{s=1}^{n}a_{s}-\sum_{s=1}^{m}c_{s}|}{\sum_{s=1}^{p}d_{s}}\right)}_{x}\right)}_{t}$ (1c)
In step 1, a trained model is used to classify all the samples. The ratio of the biased samples to the total samples predicted is determined. The tp, fp, tn, and fn are values of the true positives, false positives, true negatives, and false negatives, respectively. Since there’s hardly a perfect classifier, the positive error rate is usually reported. False positives are known to exist in similar classification systems like spam detection and automatic hate speech detection (Heron, 2009; Feng et al., 2018; Adewumi et al., 2023b).
Step 2 is similar to term frequency-inverse document frequency (TF-IDF) in that it is based on term frequency (Salton & Buckley, 1988; Ramos et al., 2003), Biased samples from step 1 are evaluated token-wise along all possible bias axes, using all the lexica of sensitive terms. An axis is a domain such as gender or race. Tables 2 and 3 provide the lexica sizes. For English and Swedish, we use the same lexica released by Alkhaled et al. (2023) and Adewumi et al. (2023c), respectively. For the other 3 languages, we create new lexica of terms (e.g. she & her) associated with specific gender or stereotypes from public sources.444fluentu.com/blog/italian/italian-nouns, en.wiktionary.org/wiki/Category:Italian_offensive_terms, Dutch_profanity, Category:German_ethnic_slurs Some of the terms in the lexica were selected from the sources based on the topmost available. These may also be expanded as needed, since bias terms are known to evolve (Haemmerlie & Montgomery, 1991; Antoniak & Mimno, 2021). The non-English lexica are small because fewer terms are usually available in other languages compared to the high-resource English language and we use the same size across the languages to be able to compare performance somewhat.
Equation 1c first finds the absolute difference between the two maximum summed frequencies in the types of an axis ($|\sum_{s=1}^{n}a_{s}-\sum_{s=1}^{m}c_{s}|$), where n and m are the total terms in a sentence along an axis. For example, in the sentence ´Women!!! PERSON taught you better than that. Shame on you!’, female terms = 1 while male terms = 0. This is then divided by the summed frequencies of all the terms ($d_{s}$) in that axis ($\sum_{s=1}^{p}d_{s}$). The operation is performed for all axes ($q$) and the average taken ($\frac{1}{q}\sum_{x=1}^{q}$). It is performed for all the biased samples ($r$) and the average taken ($\frac{1}{r}\sum_{t=1}^{r}$ ).
Axis Type 1 Type 2 Type 3
Racial 84 (black) 127 (white)
Gender 76 (female) 46 (male)
Religious 180 (christian) 465 (muslim) 179 (hindu)
Table 2: English lexica sizes (Alkhaled et al., 2023). An example of racial-black is negro.
Axis Type 1 Type 2
Gender 17 (female) 17 (male)
Racial 10 (black) 10 (white)
Table 3: Italian, Dutch, German and Swedish Lexica sizes. An example of gender-female is woman.
4 Methodology
4.1 Datasets
4.1.1 MAB
comment_text label old_id id
A fitness ranking that drops each month as more and more fat people move to Colorado.
biased
6231681
167
Een fitness ranking die elke maand daalt als meer en meer dikke mensen verhuizen naar Colorado.
Ein Fitness-Ranking, das jeden Monat fällt, wie mehr und mehr fette Menschen nach Colorado zu bewegen.
Un ranking di fitness che scende ogni mese come sempre più persone grasse si spostano in Colorado.
En fitness ranking som sjunker varje månad som fler och fler feta människor flyttar till Colorado.
Women ! ! ! PERSON taught you better than that. Shame on you!
biased
6233025
628
Vrouwen ! ! ! PERSOON heeft je beter geleerd dan dat. Je moet je schamen!
Frauen!!!! PERSON lehrte Sie besser als das. Schande über Sie!
Donne ! ! ! Person ti ha insegnato meglio di così, vergognati!
Kvinnor ! ! !- Han lärde dig bättre än så. Skäms på dig!
"The article itself touches on the topic of abortion, so absolutely it’s appropriate to discuss it."
unbiased
451965
1486981
"Het artikel zelf raakt het onderwerp abortus, dus het is absoluut gepast om het te bespreken."
"Der Artikel selbst berührt das Thema Abtreibung, also ist es absolut angebracht, es zu diskutieren."
"L’articolo stesso tocca il tema dell’aborto, quindi è assolutamente opportuno discuterne."
"Artikeln själv berör ämnet abort, så absolut är det lämpligt att diskutera det."
Table 4: English, Dutch, German, Italian, and Swedish examples from the MAB dataset. "PERSON" is the anonymization of a piece of personal identifiable information (PII)in the dataset.
Set Biased Unbiased Total
Training 533,544 1,209,433 1,742,977
Validation 32,338 69,649 101,987
Test 33,470 68,541 102,011
599,352 1,347,623 1,946,975
Table 5: MAB dataset split
The Italian, Dutch and German datasets were machine-translated from MAB  [^5] The reference provides details of the annotation of the base data. (Alkhaled et al., 2023) with the Helsinki-NLP model (Tiedemann & Thottingal, 2020). Each translation took about 48 hours on one GPU. Examples from the data are provided in Table 4. Table 5 provides statistics about the datasets. For quality control (QC), we verified translation by back-translating some random samples using Google NMT. Personal identifiable information (PII) was removed from the MAB dataset using the spaCy library. The 3 datasets are used to train new bias classifiers. We also train on the original English and the Swedish.
Machine-Translation issues:
Culture-specific biases may not be represented in the MAB versions for the translated languages because the original dataset is in English. This is a limitation. However, bias is also a universal concern, such that there are examples that span across cultures. For instance, examples 1 and 2 in Table 4 are of universal concern because individuals with non-conforming bodies and women should be respected, regardless of culture or nationality. Hence, the MAB versions are relevant for bias detection, though they were translated.
4.1.2 Evaluation datasets
Ten datasets are evaluated for bias in this work. All are automatically preprocessed before evaluation, the same way the training data were preprocessed. This includes removal of IP addresses, emojis, URLs, special characters, emails, extra spaces, numbers, empty text rows, and duplicate rows. All texts are then lowercased.
We selected datasets that are available on the HuggingFace (Wolf et al., 2020) Datasets. We evaluated the first 1,000 samples of each training split due to resource constraints. For English, we evaluated the sentence column of Corpus of Linguistic Acceptability (CoLA)(Warstadt et al., 2019), the sentence column of Question-Answering Natural Language Inference (QNLI)(Wang et al., 2018), the sentennce1 column of Microsoft Research Paraphrase Corpus (MRPC)(Dolan & Brockett, 2005), the premise column of Multi-Genre Natural Language Inference (MNLI)(Williams et al., 2018), the premise column of the CommitmentBank (CB)dataset (De Marneffe et al., 2019), and the passage column of Reading Comprehension with Commonsense Reasoning Dataset (ReCoRD)(Zhang et al., 2018). For Italian, we evaluated the context column of the Stanford Question Answering Dataset (SQuAD)(Croce et al., 2018; Rajpurkar et al., 2016); for Dutch, the sentence1 column of the Semantic Textual Similarity Benchmark (STSB)(Cer et al., 2017); for German, the text column of the German News Articles Datasets 10k (GNAD10)(Schabus et al., 2017); for Swedish, the premise of the CB.
4.2 Annotation for the assumption confirmation
To verify the assumption that toxic comments contain bias, we randomly selected 200 samples from the training set of MAB-English for annotation on Slack, an online platform. The selection of 200 samples is based on an error margin of 7% and a confidence level of 95%. To ensure high-quality annotation, we use established techniques for this task: 1) the use of gold (30) samples, 2) multiple (i.e. 3) annotators, and 3) minimum qualification of undergraduate study for annotators. Each annotator was paid 25 U.S. dollars and the it took about 2 hours to complete the annotation on average. We mixed the 30 gold samples with the 200, to verify the annotation quality of each annotator, as they were required to get, at least, 16 correctly for their annotation to be accepted. The 30 gold samples are samples with unanimous agreement in the original Jigsaw or SBICv2 data, which make up the MAB. We provide inter-annotator agreement (IAA)using Jaccard similarity coefficient (intersection over union) and credibility unanimous score (CUS)(Adewumi et al., 2023a) (intersection over sample size).
4.3 Experiments
We selected two state-of-the-art (SotA)pre-trained, multilingual models for experiments to compare their macro F1 performance: mT5-small and mBERT-base. These are from the HuggingFace hub. We further report the mT5 positive error rate of predictions. The mT5-small has 300 million parameters (Xue et al., 2021) while mBERT-Base has 110 million parameters. We trained only on the MAB datasets and evaluated using only the mT5 model, the better model of the 2, as will be observed in Section 5. For the CB and ReCoRD datasets, we evaluate all samples since they contain only about 250 and 620 entries, respectively. We used wandb (Biewald, 2020) for hyper-parameter exploration, based on Bayesian optimization. For mT5, we set the maximum and minimum learning rates as 5e-5 and 2e-5 while the maximum and minimum epochs are 20 and 4, respectively. One epoch is equivalent to the ratio of the total number of samples to the batch size (i.e. the steps). We used a batch size of 8 because higher numbers easily resulted in memory challenges.
For mBERT, we set the learning rates and epochs as with mT5. However, we explore over batch sizes of 8, 16 and 32. For both models, we set the maximum input sequence length to 512. Training took, on average, about 7.3 hours per language per epoch for mBERT while it was 6 hours for mT5. For all the experiments, we limit the run counts to 2 per language because of the long training time each takes on average. The average scores of the results are reported. The saved models with the lowest losses were used to evaluate the datasets. All the experiments were performed on two shared Nvidia DGX-1 machines that run Ubuntu 20.04 and 18.04. One machine has 8 x 40GB A100 GPUs while the other has 8 x 32GB V100 GPUs.
The lexica baseline, compared in experiments, is similar to the equation of the second step in bipol. It does not consider bias semantically and uses term frequencies, similarly to TF-IDF. It uses the same lexica as bipol. Its maximum and minimum values are 1 and 0, respectively.
5 Results and Discussion
From Table 6, we observe that all mT5 results are better than those of mBERT across the languages. The two-sample t-test of the difference of means between all the corresponding mT5 and mBERT scores have p values $<$ 0.0001 for alpha of 0.05, showing the results are statistically significant. It appears better hyper-parameter search may be required for the mBERT model to converge and achieve better performance. The best macro F1 result is for English mT5 at 0.787. This is not surprising, as English has the largest amount of training data for the pre-trained mT5 model (Xue et al., 2021). This occurred at the learning rate of 2.9e-5 and step 1,068,041.
macro F1 $\uparrow$ (s.d.) mT5 error $\downarrow$
MAB version mBERT mT5 fp/(fp +tp)
English 0.418 (0.01) 0.787 (0) 0.261
Italian 0.429 (0) 0.768 (0) 0.283
Dutch 0.419 (0.01) 0.768 (0) 0.269
German 0.418 (0.01) 0.769 (0) 0.261
Swedish 0.418 (0.01) 0.768 (0) 0.274
Table 6: Average F1 scores on the validation sets.
bipol scores $\downarrow$ (s.d.)
English $\mathit{b_{c}}$ $\mathit{b_{s}}$ bipol (b) baseline $\downarrow$
CB 0.096 0.875 0.084 (0) 0.88
CoLA 0.101 0.943 0.095 (0) 0.958
ReCoRD 0.094 0.852 0.025 (0) 0.829
MRPC 0.048 0.944 0.045 (0) 0.957
MNLI 0.063 0.833 0.053 (0) 0.965
QNLI 0.03 0.933 0.028 (0) 0.945
Italian
SQuAD 0.014 0 0.014 (0) 0.989
Dutch
STSB 0.435 0.992 0.432 (0) 0.987
German
GNAD10 0.049 0.502 0.025 (0) 1
Swedish
CB 0.08 0.938 0.075 (0) 0.97
Table 7: Average bipol & lexica baseline scores.
Figures 1 and 2 depict the validation sets macro F1 and loss line graphs for the 2 runs for the 5 languages, respectively. From Table 7, we observe that all the evaluated datasets have biases, though seemingly little (but important) when compared to the maximum of 1. We say important because many of the datasets contain small number of samples yet they can be detected. Furthermore, a low value does not necessarily diminish the weight of the effect of bias in society or the data but we leave the discussion about what amount should be tolerated open for the NLP community. Our recommendation is to have a bias score as close to zero as possible. On the other hand, the lexica baseline appears overly confident of much more bias, which is incorrect because the method fails to exclude unbiased text in its evaluation, which is a shortcoming of methods based solely on it. The Dutch STSB is higher than the other bipol scores because of the higher bipol classifier component score of 0.435, which may be because of the nature of the dataset.
Figure 1: Macro F1 of the validation set for the 5 languages, as generated by wandb.
Figure 2: Loss on the validation set for the 5 languages, as generated by wandb.
Figure 3: Error matrix of mT5 on MAB-English
5.1 Error analysis & qualitative results
According to the error matrix in Figure 3, the mT5 model is better at correctly predicting unbiased samples. This is because of the higher unbiased samples in the training data of MAB. In Table 8, the first example for the English CB contains a stereotypical statement "men are naturally right and it is the role of women to follow their lead", leading to the correct biased prediction by the model. Similarly, this correct prediction is made in the Swedish CB. We notice over-generalization (May et al., 2019; Nadeem et al., 2021) in the correct examples for the CoLA predictions, where "every" is used. The table also shows some incorrect predictions.
comment_text prediction csv id
Correct predictions from the English CB
nicky approached her with the assumption that men are naturally right and it is the role of women to follow their lead. constance whose confidence was growing daily was not prepared to give in to nickys wishes merely because of his sex. if she felt he was right then she agreed with him.
biased
85
b: thats true. a: so. b: uh the other argument is that the death penalty is a deterrent and i really dont uh agree with that. i dont think anyone who would commit uh a crime that would get them the death penalty would stop at the moment and say well i was about to kill and dismember this person but oh if they catch me theyre going to kill me so i better not do it. i just dont think uh that it works that way. a: yeah. i dont think its done.
biased
133
Incorrect prediction
b: yeah and the ground will filter some of it but not all of it. a: no not when you figure i didnt realize one cow produces that much manure
biased
137
Correct predictions from the English CoLA
if you give him enough opportunity every senator will succumb to corruption.
biased
266
every senator becomes more corrupt the more lobbyists he talks to.
biased
277
Incorrect prediction
bill squeezed the puppet through the hole.
biased
82
Table 8: Qualitative examples of apparently correct and incorrect predictions in some of the English datasets, based on the mT5 model.
5.2 Consistent prediction with perturbation
An interesting property of relative consistency that we observed with the model predictions, as demonstrated with the CoLA dataset, is that when sentences are perturbed, the model mostly maintains its predictions, as long as the grounds for prediction (in this case - over-generalization) remain the same. The perturbations are inherent in the CoLA dataset itself, as the dataset is designed that way. Some examples are provided in Table 9, where 6 out of 8 are correctly predicted. This property is repeated consistently in other examples not shown here.
comment_text prediction csv id
if you give him enough opportunity every senator will succumb to corruption.
biased
266
you give him enough opportunity and every senator will succumb to corruption.
biased
267
we gave him enough opportunity and sure enough every senator succumbed to corruption.
unbiased
268
if you give any senator enough opportunity he will succumb to corruption.
biased
269
you give any senator enough opportunity and he will succumb to corruption.
biased
270
you give every senator enough opportunity and he will succumb to corruption.
biased
271
we gave any senator enough opportunity and sure enough he succumbed to corruption.
biased
272
we gave every senator enough opportunity and sure enough he succumbed to corruption.
unbiased
273
Table 9: Mostly consistent correct prediction with perturbation in the CoLA dataset.
5.3 Explainability by graphs
We show explainability by visualization using graphs. Bipol produces a dictionary of lists for every evaluation and we represent 2 in top-5 frequent terms bar graph. Many of the datasets display overall male bias. As examples, figures 4 and 5 show that the GNAD10 and STSB datasets have overall male and female bias, respectively.
Figure 4: Top 5 frequent terms in the GNAD10 dataset (paired terms are only for comparison).
Figure 5: Top 5 frequent terms in the STSB dataset (paired terms are only for comparison).
5.4 Assumption confirmation through annotation
The results of the annotation of the 200 MAB samples reveal that toxic comments do contain bias. This is shown in Figure 6. The Jaccard similarity coefficient and CUS of IAA are 0.261666Not to be interpreted using Kappa for 2 annotators on 2 classes. Ours involved 3 annotators and 0.515, respectively, given that over 50% is the intersection of unanimous decision.
Figure 6: Annotation confirms assumption about toxic comments.
6 Conclusion
The findings of this work show that bias besets Natural Language Processing (NLP)datasets regardless of language, including benchmark datasets on the GLUE/SuperGLUE leaderboards. We introduced MAB datasets in 3 languages for training models in bias detection. Each has about 2 million labeled samples. We also contribute lexica of bias terms for the languages. In addition, we verified the assumption that toxic comments contain bias. It may be impossible to completely remove bias from data or models, since they reflect the real world, but resources for estimating bias can provide insight into mitigation strategies for reducing bias. Future work may explore ways of minimising false positives in classifiers to make them more effective. One may also explore how this work scales to other languages or how multilignual models compare to language-specific monolingual models or large language models (LLM s). Regarding culture-specific biases in datasets, one solution will be to collect data from the specific cultures/nationalities to capture these biases.
Ethics statement & limitation
The authors took care while providing examples of data samples, despite some containing stereotypes or toxic content. The classifiers trained for estimating the biases in this work are limited in effectiveness, as shown in the results, hence a result of 0 on any dataset does not necessarily indicate a bias-free dataset. The original MAB was annotated by humans who may have their personal biases based on cultural or demographic backgrounds. This means the final annotations may not be seen as absolute ground truth of social biases.
Acknowledgments
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation and counterpart funding from Luleå University of Technology (LTU). The authors also thank Björn Backe for his insightful comments during the writing of this paper.